### v1

# Geometry-Guided Ray Augmentation for Neural Surface Reconstruction with Sparse Views

Abstract
In this paper, we propose a novel method for 3D scene and object reconstruction from sparse multi-view images. Different from previous methods that leverage extra information such as depth or generalizable features across scenes, our approach leverages the scene properties embedded in the multi-view inputs to create precise pseudo-labels for optimization without any prior training. Specifically, we introduce a geometry-guided approach that improves surface reconstruction accuracy from sparse views by leveraging spherical harmonics to predict the novel radiance while holistically considering all color observations for a point in the scene. Also, our pipeline exploits proxy geometry and correctly handles the occlusion in generating the pseudo-labels of radiance, which previous image-warping methods fail to avoid. Our method, dubbed Ray Augmentation (RayAug), achieves superior results on DTU and Blender datasets without requiring prior training, demonstrating its effectiveness in addressing the problem of sparse view reconstruction. Our pipeline is flexible and can be integrated into other implicit neural reconstruction methods for sparse views.
1 Introduction
Figure 1: With a limited number of posed images (left), our proposed RayAug could produce high-quality novel view synthesis and accurate surface reconstruction when training on clean images (right). We reconstruct the meshes using the marching cubes algorithm [17].
Reconstructing 3D geometry from multi-view inputs is a fundamental and challenging problem in computer graphics and computer vision. Traditional multi-view stereo (MVS) methods primarily rely on correspondence matching between input images, followed by fusing predicted depth maps into point clouds for the final meshing process. However, these techniques require hand-crafted heuristics, which may struggle to find accurate matches that are crucial for producing meaningful dense surfaces when dealing with sparse input images.
Recent advances in neural implicit representations have shown promising results for 3D scene representation. The groundbreaking neural radiance fields (NeRF) [19] leverages a multi-layer perceptron (MLP) to predict per-point radiance and density and uses differentiable volume rendering to synthesize novel views. Neural surface reconstruction [32, 35, 36] methods also leverage radiance fields based techniques to supervise view synthesis for surface reconstruction. However, like traditional approaches, they also require dense images to reconstruct intricate geometry, which limits their applicability.
To address the sparse view problem, some methods have been proposed that enhance supervision during training. These approaches primarily use multi-view consistency to create pseudo labels for unseen rays by re-projection [39] or image warping [33] 3D points to unseen viewpoints. While these techniques improve results in sparse view settings, their augmentation strategies face two key issues: (1) they do not guarantee that the reprojected or warped points are visible at the augmented camera poses, and (2) they simply assign the color of the augmented rays based on the source view, leading to inaccurate estimations.
In this paper, we address these limitations and develop a method to generate more precise pseudo-labels for training radiance fields. We empirically observe that the rough geometry in the early stage of the implicit surface representation training provides valuable information with unbiased surface estimation and regularization. By combining this insight with physical priors, we propose RayAug, a novel geometry-guided approach for neural reconstruction from sparse views.
Specifically, we shoot rays from 3D points on the surface of the rough geometry to unseen viewpoints, ensuring their observability by ray marching the signed-distance field. Rays that hit the surface are discarded. Additionally, since a 3D point can be seen from multiple input viewpoints, we aggregate all the color observations by fitting spherical harmonics (SH) for each spatial point. This approach aligns with the distribution of natural radiance, providing better extrapolation results than linear methods. The color of each augmented ray from the same 3D point is then obtained by querying the fitted spherical harmonics function for each point. Finally, we perform another visibility check to avoid biased estimation of color caused by occlusion. We evaluate our method through novel view synthesis and mesh reconstruction on DTU dataset [10] and Blender dataset [19], demonstrating significant improvements compared to baseline methods.
In summary, our contributions are as follows:
  • We exploit spherical harmonics to holistically consider all color observations for a 3D point, enabling pseudo-label supervisions consistent with the natural distribution of radiance.
  • We propose to use rough geometry to avoid occlusion in the ray augmentation process, which avoids the confound of point radiance and reconstructs more accurate surfaces from sparse views.
  • Our method works as a plug-and-play method and can be incorporated with other techniques for sparse implicit neural surface reconstruction. Our method achieves superior results on sparse view reconstruction without the need for extra data and prior training.
2 Related Work
2.1 Multi-view Reconstruction
The traditional multi-view stereo (MVS) pipeline first estimates per-view depth maps by incorporating and matching images. Then the depth maps are filtered and fused to obtain a global representation i.e. point clouds [7, 27]. 3D surfaces can thus be reconstructed by applying meshing algorithms such as Poisson reconstruction [13] or Delaunay triangulation [15] on point clouds. Later, learning-based approaches provide more accurate depth estimation by exploiting massive data and achieving better performance on MVS tasks [34, 8]. These methods produce promising results with dense captured images [26], but rely heavily on the quality of correspondence matching. Therefore, they are sensitive to outliers, noises, and textureless regions, leading to incomplete and noisy reconstructions.
2.2 Neural Surface Reconstruction
Recently, neural implicit representations have demonstrated superior effectiveness in various tasks such as 3D object [23, 18, 3] and scene representation [19, 29, 28], novel view synthesis [19, 1, 31] and multi-view 3D reconstruction [35, 21, 32, 22]. Among them, NeRF and its variants [19, 1] demonstrated state-of-the-art performance for novel view synthesis with differentiable volumetric rendering. However, they require dense images for input and it is hard to extract high-quality geometry because the predicted density field lacks sufficient constraints. Other works utilizing surface rendering [21] produce more accurate modeling but require additional constraints such as ground truth masks or depth priors for supervision [38, 35]. Several later studies combine the advantages of surface and volume representation, generating satisfying surfaces with only image inputs [32, 36, 22]. For example, NeuS [32] combined volume rendering with signed distance functions to circumvent the problem of insufficient surface constraint in the original NeRF. The sampling efficiency was further improved to allow for reconstructing in the wild scenes [30]. UNISURF [22] also unifies surface and volume rendering but with occupancy fields. Similar to them, our work also only requires multi-view images for 3D reconstruction.
2.3 Implicit Representation with Sparse Views
Several works specifically deal with implicit representations given sparse views as it often happens in real-world applications. A common way to alleviate the requirements of dense inputs is to learn a generalized network from multiple scenes. These methods often contain a learned image encoder to encode inputs and then fuse multi-view features to assist the volume rendering process [37, 2, 11, 25, 16]. However, they require additional datasets for training and have potential domain generalization issues.
Figure 2: Method Overview. We exploit the coarse geometry in radiance field training to guide its augmentation with sparse inputs. (left) For a surface point $\mathbf{v}$, we aggregate the color observations from all inputs to fit a spherical harmonics expansion, thus the pseudo labels for all the augmented rays passing through $\mathbf{v}$ can be obtained through querying the SH, (right) when generating augmented rays, we check their visibility and exclude those cannot be actually observed.
Another direction of work avoids pre-training through novel supervision in different aspects. DS-NeRF [5] and StructNeRF [4] supervise NeRF with sparse depth from structure-from-motion. The latter further exploits multi-view image and plane constraints for indoor scenes. DietNeRF [9] enforces semantic consistency of rendered patches with CLIP [24], which only provides inadequate high-level information. RegNeRF [14] regularizes the geometry and appearance of patches rendered from unobserved viewpoints, reducing the floating artifacts normally presented for limited input views. InfoNeRF [14] proposes regularizers for NeRF instead from the ray entropy perspective and can be flexibly plugged into other NeRF models. RayAug also has no pre-training procedure, but we differ from those works in that instead of introducing regularization or across scenes prior, we try to fit a radiance distribution of each surface point and use this as the augmented information. More specifically, we assume that the radiance distribution of a single point is a harmonic function and can be decomposed into a spherical harmonic expansion. Following this idea, we design a physically grounded pipeline to create precise pseudo labels for radiance supervision.
3 Preliminaries: NeuS
NeuS [32] represents a 3D scene as two continuous functions both parametrized by a multi-layer perception (MLP). The first one is a signed-distance function that takes as input 3D position $\mathbf{x}=(x,y,z)$ and predicts its signed distance: $f_{1}(\mathbf{x})\rightarrow s$, and the other takes both $\mathbf{x}$ and its viewing direction $\mathbf{d}$ and predicts its view-dependent color: $f_{2}(\mathbf{x},\mathbf{d})\rightarrow\mathbf{c}$.
Similar to NeRF [19], the color of a pixel in NeuS [32] only depends on the radiance along a ray with no other lighting factors. For a camera ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ with center $\mathbf{o}$ and direction $\mathbf{d}$, its color can be derived with volume rendering [12]. In practice, NeuS samples $N$ points along $\mathbf{r}(t)$ with $t=t_{1},...t_{N},t_{i}<t_{i+1}$ and approximate the per-pixel color with the following:
$\hat{\mathbf{C}}(\mathbf{r})$ $=\sum_{i=1}^{N}T_{i}\alpha_{i}\mathbf{c}_{i},$ (1)
$T_{i}$ $=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (2)
where $\alpha_{i}$ is the discrete opacity value that is a function of the probability density function of $f_{1}$, and $T_{i}$ is the accumulated transmittance which indicates the probability that a ray traverses from $t_{1}$ to $t_{i}$ without hitting any particle. NeuS exploits volume rendering to minimize the mean-squared error (MSE) between the predicted ray color $\hat{\mathbf{C}}(\mathbf{r})$ and the ground truth $\mathbf{C}(\mathbf{r})$:
$\mathcal{L}_{\mathrm{MSE}}=\sum_{\mathbf{r}}\|\hat{\mathbf{C}}(\mathbf{r})-\mathbf{C}(\mathbf{r})\|_{2}^{2}.$ (3)
Besides color loss, it also leverages an L1 regularization and a mask loss if available. With successful minimization, the formulation of NeuS enables the network to learn a signed-distance function that represents an accurately reconstructed surface.
4 Methods
When limited images and corresponding training rays are given, from 3, only a small proportion of the 3D space has supervision, while the results of the rest space purely rely on the interpolation ability of the network. Therefore, the main goal of our paper is to augment the supervision signal of Implicit Neural Representation (INR) by inferring rays from unseen viewpoints based on limited inputs and the reconstructed INR in an accurate way. Particularly, we found that model the problem of inferring unseen radiance as fitting the radiance maps for spatial points, and the radiance map of each point could be estimated from the known radiance of that point based on our prior radiance map function.
In the following of this section, we introduce the details of our method. Our radiance augmentation algorithm is composed of three steps: Surface Ray Casting in 4.1, Visibility Checking in 4.2, and Radiance Map Estimation in 4.3.
4.1 Surface Ray Casting
Same to NeRF, the input to our method is a set of posed images $\{I_{i}\}_{i=1}^{N}$, where $N$ is the number of images. The original training set is denoted by =$\mathcal{R}=\{(\mathbf{r},\mathbf{c})\}$, in which $\mathbf{r}$ and $\mathbf{c}$ are the ray directions and ray color. Since our method relies on the rough scene geometry, we first train a radiance field for a few thousand steps and construct the mesh with marching cubes [17]. The vertices of this rough mesh are regarded as points lying on the surface. Then starting from these surface points $\mathbf{v}$, we shoot a number of rays $\mathbf{r}_{rand}\in\mathcal{R}$ in random directions. We denote the ray casted as ~$\tilde{\mathcal{R}}_{\mathbf{v}}$. We ensure that the angle between new rays and the surface normal of $\mathbf{v}$ is less than $\frac{\pi}{2}$, so these candidates for augmented rays are confident.
4.2 Visibility Check
However, some rays in ~$\tilde{\mathcal{R}}_{\mathbf{v}}$ might not actually be observed from a valid viewport since they will hit the mesh and get occluded. Figure 2 shows an example of this case. To filter out those invalid rays, we include an additional visibility check step to exclude those invalid rays. Specifically, for each ray $\mathbf{r}_{aug}$, we march from the starting point $\mathbf{v}$ according to the scene signed-distance field (SDF) and terminate when the surface is reached or achieve the maximum number of steps. Since the ray marching method is efficient, we integrate it directly into the training pipeline without any offline computation.
Algorithm 1 Ray Marching for Visibility Check
Input: pos (starting point), $sdf$ (signed-distance function)
Output: visible or not
1:step := 0
2:d := $sdf(\text{pos})$
3:while d $>\epsilon$ and step $<$ max_steps do
4:     pos = pos + dir * d
5:     d = $sdf(\text{pos})$
6:     step = step + 1
7:end while
8:if step $<$ max_steps then
9:     return true // visible
10:else
11:     return false // invisible
12:end if
4.3 Radiance Map
After the visibility check, rays in ~$\tilde{\mathcal{R}}_{\mathbf{v}}$ are ensured to be observed from at least one camera viewpoint. The next question is what’s the proper radiance value of rays in ~$\tilde{\mathcal{R}}_{\mathbf{v}}$? In previous ray augmentation methods [39], a surface ray has only one known radiance value, since the source view visibility check is not performed, thus they simply assign a color from one input view, which neglects the view-dependent effect and leads to inconsistent pseudo label. However, as shown in the previous discussion, a point in the scene surface could be seen from multiple views and we actually have more information regarding the radiance of surface point $\mathbf{v}$. Thus, a main contribution of our method illustrated as follows is that we utilize all reasonably credible information to reconstruct the radiance distribution of surface point $\mathbf{v}$ to infer the radiance of novel rays. Actually, not only did we fit the radiance distribution but also we achieved extrapolation using spherical harmonics expansion fitting.
Our method handles the inference of radiance in two situations. Firstly in the simpler case, if the point can only be seen from less than $N_{v}$ views ($N_{v}$ is empirically 10 in our experiment), we think that we don’t have enough information to reconstruct the radiance distribution. Thus, the radiance value of a randomly chosen view is assigned to all new rays origin from point $\mathbf{v}$. This is basically in accordance with the previous method.
If the point $\mathbf{v}$ under processing can be viewed from more than $N_{v}$ views, then in this case we try to infer a distribution of radiance over all viewing directions. Our solution to this problem is to fit a spherical harmonic expansion, which is widely used in precompute radiance transfer and environment maps in computer graphics. We represent the radiance distribution using spherical harmonics:
$\mathbf{c_{v}}(\mathbf{d})=\sum_{\ell=0}^{\ell_{\mathrm{max}}}\sum_{m=-\ell}^{\ell}k_{\ell}^{m}Y_{\ell}^{m}(\mathbf{d}),$ (4)
where $\mathbf{c_{v}}(\mathbf{d})$ is the radiance of point $\mathbf{v}$ viewing from $\mathbf{d}$. Namely, for vertice v, we back-project it to $\{I_{i}\}_{i=1}^{N}$ and filter views that according to visibility (use the same visibility check function as in 4.2). The color observations for v can thus be represented by:
$\textbf{S}_{\mathbf{v}}=\{(\mathbf{c}_{j},\mathbf{d}_{j})\}=\{(\textbf{c}_{\textbf{v},j},\mathrm{Norm}(\textbf{o}_{j}-\textbf{v}))\}_{j=1}^{N_{\mathrm{vis}}},$ (5)
where $N_{\mathrm{vis}}$ is the number of views from where v can be seen and $\textbf{o}_{j}$ is the corresponding camera center. These radiance sample points are used to fit a radiance map by computing the coefficients for the SH expansion via least squares fitting:
$\mathop{\arg\min}\limits_{k_{\ell}^{m}}\>\sum_{j=1}^{N_{vis}}|\mathbf{c_{v}}(\mathbf{d}_{j})-\mathbf{c}_{j}|^{2},$ (6)
where $k_{\ell}^{m}\in\mathbb{R}^{3}$ is a set of 3 coefficients for RGB components. Then, for each ray in ~$\tilde{\mathcal{R}}_{\mathbf{v}}$, color is determined by querying the SH functions $Y_{\ell}^{m}$ given its viewing direction $\mathbf{d}$. Using such an SH expansion to fit the radiance distribution can provide extrapolation for novel view direction.
Figure 3: Visualization of radiance fitting. We project the radiance value of points distributed on the upper hemisphere into this circle to get a visualization of the fitted radiance distribution. The first row of images depicts the SH fitting results, while the second row showcases the outcomes of interpolation. It reveals that the SH fitting provides more informative radiance prediction results on the top line, while preserving the smoothness of distribution.
To summarize, our method extends the training rays of NeRF considering scene geometry. For each surface vertice v, we first construct ~$\tilde{\mathcal{R}}_{\mathbf{v}}$ by shooting random ray directions from v and filter invisible ones through visibility check. Since v can be observed at multiple input viewpoints, we model it as an SH which thus can be queried at any augmented ray direction going through v.
Figure 4: Qualitative results of novel view synthesis on DTU dataset [10] and Blender dataset [19]. We train the model using 30% images for real scenes in DTU dataset, and 6% for synthetic scenes in Blender dataset.
5 Experiments
Figure 5: Qualitative comparison of the reconstructed depth and foreground mask. The upper row is the result without novel radiance prediction. The lower row is the result using our method.
Known/Unknown 8:2 7:3 6:4 5:5
Interpolation 0.0571 0.0633 0.0101 0.1540
SH Fitting 0.0826 0.0770 0.0897 0.1056
Table 1: MSE in the novel views between interpolation and SH fitting with different ratios between known rays and unknown rays.
Chair
Ficus
Lego
NeuS
MVSNeRF
SparseNeuS
Ours
Figure 6: Qualitative results of mesh reconstruction on Blender dataset [19].
Ratio (Used/All) $\alpha=0.2$ $\alpha=0.3$ $\alpha=0.4$
[…]
MVSNeRF 24.65 7.53 24.77 5.67 25.85 5.23
Depth-supervised NeRF 23.41 5.41 22.07 6.41 23.97 6.64
SparseNeuS-ft 24.90 3.70 25.35 3.21 26.82 3.04
NeuS 22.85 5.45 23.09 4.21 23.41 3.41
RayAug (Ours) 26.81 3.58 27.08 3.17 27.59 3.12
Table 2: Quantitative comparison against the other state-of-the-art methods on DTU dataset [10]. We evaluate both novel view synthesis (PSNR) and reconstructed mesh quality (Chamfer Distance) on 20 to 40 percent of the training images. Each training image is randomly chosen.
5.1 Novel Radiance Predicting from Spherical Harmonics
Firstly, we conduct an experiment to validate the effective method for generating radiance from unseen views. For predicting the novel radiance, we compare our method with a naive baseline: spherical linear interpolation, which interpolates the radiance of the unseen view from the nearest two seen views in the spherical coordinate system. Specifically, we implement the linear interpolation of radiance using the geodesic distance as the weight:
$\mathbf{\hat{c}}=w\mathbf{c}_{0}+(1-w)\mathbf{c}_{1},$ (7)
$w=\frac{d(\mathbf{\hat{v}},\mathbf{v}_{1})}{d(\mathbf{v}_{1},\mathbf{\hat{v}})+d(\mathbf{\hat{v}},\mathbf{v}_{2})},$ (8)
$d(\mathbf{v}_{1},\mathbf{v}_{2})=\arctan\frac{\left|\mathbf{v}_{1}\times\mathbf{v}_{2}\right|}{\mathbf{v}_{1}\cdot\mathbf{v}_{2}},$ (9)
while $\mathbf{\hat{c}},\mathbf{\hat{v}}$ are the radiance and view direction vector of novel view, $\{\mathbf{c}_{i},\mathbf{v}_{i}|i\in{1,2}\}$ are these of the nearest two views, respectively. Intuitively, the spherical harmonics fitting method can provide extrapolation of radiance compared to the linear interpolation method while preserving the smoothing variation under the prior of harmonic function.
We generate a set of around $10,000$ points with all of their visible colors from the DTU dataset for evaluating the two methods for radiance prediction. The radiance from different views of each point is divided into two training (known) and testing (unknown) views with ratios from $8:2$ to $5:5$, in order to evaluate the performance under the different number of novel views. We report the average normalized Mean Square Error (MSE) in the novel views in Table 1. The average MSE for SH fitting is less than that of interpolation, especially when the number of novel views exceeds the number of known views. A visualization case is shown in Figure 3.
Methods Images Chair Drums Ficus Hot Dog Lego Materials Mic Ship Average
NeuS Ours 6 9
21.41 13.70 20.83 19.74 18.36 14.98 21.72 16.47 20.28
24.59 16.91 23.57 27.73 24.17 18.70 25.97 19.92
22.28 15.69 21.71 22.89 19.96 15.95 23.62 18.62 21.59
27.59 19.37 24.09 26.61 23.10 18.36 27.48 21.77
Table 3: Quantitative comparison on Blender dataset [19]. We compare our method with NeuS [32], and reconstruct 8 scenes using 6 or 9 images. The table shows the PSNR on each reconstructed scene.
5.2 Implementation Details
5.2.1 Dataset
We evaluate our method on the DTU dataset [10] and reconstruct 6 scenes in the dataset. These scenes contain various materials, textures, and complex geometry shapes such as holes, multiple objects, and non-convex parts of objects. Originally, each scene contains 49 or 64 images, and the resolution for training and testing is 1600 $\times$ 1200. We also reconstruct 8 scenes from the Blender dataset [19] using our method. Each scene in the Blender dataset contains 100 training images, and scenes are generated from a physical-based rendering engine and include non-Lambertian effects such as reflections. We leverage the mask provided for training and testing, following the same setting used in [16, 32].
5.2.2 Implementation Details
In general, we implement our ray augmentation strategy based on NeuS [32]. Previous work [20] has shown that the input encoding scheme is important for reconstructing an implicit representation. Thus, we compare and choose the frequency encoding [31] for the position input, and sphere harmonics function encoding [6] for the direction. We find that by replacing the frequency encoding with SH function encoding for direct input, the rendered color can converge faster.
We train our model using the Adam optimizer with an initial learning rate of $41e-$ 4 and a learning rate warm-up strategy. We choose a batch size of 2048 and train our model for 150k steps, and each scene is trained on one NVIDIA A100 GPU. To extract mesh for ray augmentation, we run marching cubes for a spatial resolution of $64^{3}$ to get the vertices. The same algorithm is used for extracting mesh when we evaluate the reconstruction result.
5.2.3 Number of Images
The quality of reconstruction with respect to the number of images given is the main concern of this paper, especially when given a few images. For fair comparison across scenes, we do not directly extract a fixed number of images for each scene like previous paper but leverage the ratio $\alpha$ equal to the number of images used and the number of all images provided in the dataset. To evaluate the performance of view synthesis, we randomly choose 70% as the training set and the rest as the testing test. Among the training images, we further choose $\alpha=\{20\%,30\%,40\%\}$ for both image rendering and surface reconstruction.
5.3 Experiment Results
5.3.1 Quantitative
We conduct quantitative comparisons on image quality and mesh quality with NeuS [32], Depth-supervised NeRF [5], SparseNeuS [16] and MVSNeRF [2] on the DTU dataset. Additionally, we compare our method with NeuS for novel view synthesis task on the Blender dataset. For the comparison, we exploit the official public code of the aforementioned methods and employ Peak Signal-to-Noise Ratio (PSNR) and Chamfer Distance (CD) for measuring image quality and mesh quality, respectively.
In Table 2, our method demonstrates significant improvements in the quality of rendered images for all $\alpha$ compared to NeuS [32]. We also train and fine-tune SparseNeuS [16] following the settings in their paper, and the Chamfer Distance of the reconstructed mesh is better decent since they leverage cross-scene prior knowledge for training, and their method achieves a smaller distance compared to our method when using 40% of the images for training. However, in all other cases, our method reconstructs the most precise mesh as evaluated out of other tested methods. It is worth highlighting that even with a limited number of images, our strategy yields a significant improvement in mesh quality. Furthermore, despite MVSNeRF [2] leveraging deep features to enable generalizable across-scene training, our method still surpasses MVSNeRF in terms of view synthesis quality, regardless of the number of images used. The mesh quality of MVSNeRF drops down since the surface estimation is biased. Note that we compare the raw mesh obtained by marching cubes without any further cleaning to the ground truth mesh.
ScanID w/o vis check w/ vis check
scan24 12.00 15.92
scan37 12.74 13.87
scan40 11.16 15.67
scan55 13.52 17.10
scan63 14.94 23.35
scan65 14.10 19.52
scan69 12.55 14.94
scan83 18.75 24.10
scan97 13.02 17.55
scan106 11.62 13.62
scan114 13.77 18.52
scan118 14.86 19.57
scan122 14.99 18.57
mean 13.69 17.87
Table 4: Ablation Study for the effectiveness of the visibility check. We compare the view synthesis results of enabling visibility and disabling visibility when reconstructing 13 scenes from DTU dataset [10] using 3 or 5 images for each scene.
Table 3 presents the performance of RayAug on the Blender dataset which outperforms NeuS by a considerable margin. Compared with NeuS under different numbers of images, our method obtains better results on all 8 scenes, demonstrating the effectiveness of our proposed method.
5.3.2 Qualitative
We also present qualitative comparisons of our method on the DTU dataset. Figure 4 shows the results for novel view synthesis, and our rendered images preserve more details compared to other methods. The rendered images using our method are more natural due to the more radiance supervision signal predicted from unseen views. As we can see from the last row, the cymbals of the drums are brighter, which is more identical to the ground truth when using our method. The results show that our method can alleviate the floaters in the reconstructed scenes. In Figure 5, our method demonstrates satisfactory performance in reconstructing the accurate depth range. The mesh reconstruction results are also provided in Figure 6.
5.3.3 Ablation Study
We investigate the effectiveness of the visibility check and its impact on the view synthesis results of ray augmentation. Table 4 shows the results of view synthesis with and without visibility check for each reconstructed scene. We train 13 scenes from the DTU dataset and set $\alpha=0.1$, which is 3 or 5 images for each scene, depending on the original number of images. The results indicate that the visibility check plays a crucial role in the ray augmentation process since augmenting rays with occluded rays would lead to wrong estimations of novel radiance, thus significantly decreasing the performance of view synthesis.
6 Conclusion
Figure 7: Unsuccessful cases in our reconstruction. It reveals the mesh at the back of the objects have several rough regions due to lack of supervisions.
In this paper, we introduce RayAug, a novel approach to address the challenge of sparse view reconstruction. Our method enhances the training of radiance fields by employing a geometry-guided strategy: checking the visibility of augmented ray directions using coarse geometry on hand. This Visibility check prevents the introduction of incorrect rays that penetrate known surfaces, leading to noisy labels, and ensures a more physically accurate augmentation process. Moreover, the radiance map incorporates a prior of natural radiance distribution, composed of a series of harmonic functions, providing confident novel radiance and proving to be more accurate than merely assigning the color from a single image. Furthermore, RayAug can be easily integrated into any surface reconstruction technique based on volume rendering. Comprehensive experiments demonstrate our method outperforms existing methods. We believe that our techniques have the potential to become a standard tool for data augmentation in implicit representation reconstruction.
However, our work also has some limitations. Firstly, while we exploit ray augmentations to tackle the sparse view problem, regions that fail the visibility check cannot be augmented, such as the back of an object (See Figure 7). Future work could consider incorporating priors by training on a larger dataset. Additionally, in the extreme case of a single view, our spherical harmonics fitting would not be useful since there is only one observation for each input pixel. Exploring alternative solutions for such cases remains an exciting avenue for future research.